OpenStudy (anonymous):

Given the proportion a/b = 6/20 , what ratio completes the equivalent proportion a/b = / ?

OpenStudy (anonymous):

20/b a/20 6/20 b/20

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

Can you help me again @Jim766 ?

OpenStudy (jim766):

are you sure you copied the answer choices correctly? it is not making sense...

OpenStudy (anonymous):

Yeah, they're suppose to be fractions.

OpenStudy (jim766):

\[\frac{ 6 }{ 20 }=\frac{ 3 }{ 10 }\]

OpenStudy (jim766):

in that case 6/20 = 6/20 is the only one that is true